# GroupMamba: Parameter-Efficient and Accurate Group Visual State Space Model

### v1

@@ Abstract
Recent advancements in state-space models (SSMs) have showcased effective performance in modeling long-range dependencies with subquadratic complexity. However, pure SSM-based models still face challenges related to stability and achieving optimal performance on computer vision tasks. Our paper addresses the challenges of scaling SSM-based models for computer vision, particularly the instability and inefficiency of large model sizes. To address this, we introduce a Modulated Group Mamba layer which divides the input channels into four groups and applies our proposed SSM-based efficient Visual Single Selective Scanning (VSSS) block independently to each group, with each VSSS block scanning in one of the four spatial directions. The Modulated Group Mamba layer also wraps the four VSSS blocks into a channel modulation operator to improve cross-channel communication. Furthermore, we introduce a distillation-based training objective to stabilize the training of large models, leading to consistent performance gains. Our comprehensive experiments demonstrate the merits of the proposed contributions, leading to superior performance over existing methods for image classification on ImageNet-1K, object detection, instance segmentation on MS-COCO, and semantic segmentation on ADE20K. Our tiny variant with 23M parameters achieves state-of-the-art performance with a classification top-1 accuracy of 83.3% on ImageNet-1K, while being 26% efficient in terms of parameters, compared to the best existing Mamba design of same model size. Our code and models are available at: https://github.com/Amshaker/GroupMamba.
@@ 1 Introduction
However, Mamba, specifically the S6 algorithm, is known to be unstable for e.g., image classification, especially when scaled to large sizes [45]. Additionally, the Mamba model variant used in image classification generally called the VSS (Visual State Space) block, can be more efficient in terms of parameters and compute requirements based on the number of channels. The VSS block includes extensive input and output projections along with depth-wise convolutions, whose parameters and compute complexities are directly proportional to the number of channels in the input. To address this issue, we propose a Modulated Group Mamba layer which mitigates the aforementioned issues in a computation and parameter-efficient manner. The main contributions of our paper are:
@@
    We employ a distillation-based training objective to further stabilize the training of models with a large number of parameters.
@@
    We build a series of parameter-efficient generic classification models called “GroupMamba”, based on the proposed Modulated Group Mamba layer. Our tiny variant achieves 83.3% top-1 accuracy on ImageNet-1k [9] with $23M$ parameters and $4.6G$ FLOPs. Additionally, our base variant achieves top-1 accuracy of 84.5% with $57M$ parameters and $14G$ FLOPs, outperforming all recent SSM methods (see Figure 1).
Figure 1: Left: Comparison in terms of Parameters vs. Top-1 Accuracy on ImageNet-1k [9]. Our GroupMamba-B achieves superior top-1 classification accuracy while reducing parameters by 36% compared to VMamba [34]. Right: Qualitative results of GroupMamba-T on semantic segmentation (top right), and object detection and instance segmentation (bottom right). More qualitative examples are presented in Figure 3 and Figure 4
@@ 3 Method
  • Lack of Stability for Larger Models: We observe from [45] that Mamba [16] based image classification models with an MLP channel mixer are unstable when scaled to a large number of parameters. This instability can be seen in SiMBA-L (MLP) [45], which leads to sub-optimal classification results of $49\%$ accuracy. We mitigate this issue by introducing a Modulated Group Mamba design alongside a distillation objective (as presented in Section 3.4) that stabilizes the Mamba SSM training without modifying the channel mixer.
  • Efficient Improved Interaction: Given the computational impact of Mamba-based design on the number of channels, the proposed Modulated Group Mamba layer is computationally inexpensive and parameter efficient than the default Mamba and able to model both local and global information from the input tokens through multi-direction scanning. An additional Channel Affinity Modulation operator is proposed in this work to compensate for the limited channel interaction due to the grouped operation.
@@ 3.1 Preliminaries
State-Space Models: State-space models (SSMs) like S4 [17] and Mamba [16] are structured sequence architectures inspired by a combination of recurrent neural networks (RNNs) and convolutional neural networks (CNNs), with linear or near-linear scaling in sequence length. Derived from continuous systems, SSMs define and 1D function-to-function map for an input $x(t)\in\mathbb{R}^{L}\rightarrow y(t)\in\mathbb{R}^{L}$ via a hidden state $h(t)\in\mathbb{R}^{N}$. More formally, SSMs are described by the continuous time Ordinary Differential Equation (ODE) in Equation 1.
@@
where $h(t)$ is the current hidden state, $h^{\prime}(t)$ is the updated hidden state, $x(t)$ is the current input, $y(t)$ is the output, ${\mathbf{A}}\in\mathbb{R}^{N\times N}$ is SSM’s evolution matrix, and ${\mathbf{B}}\in\mathbb{R}^{N\times 1},{\mathbf{C}}\in\mathbb{R}^{N\times 1}$ are the input and output projection matrices, respectively.
Discrete State-Space Models: To allow these models to be used in sequence modeling tasks in deep learning, they need to be discretized, converting the SSM from a continuous time function-to-function map into a discrete-time sequence-to-sequence map. S4 [17] and Mamba [16] are among the discrete adaptations of the continuous system, incorporating a timescale parameter ${\mathbf{\Delta}}$ to convert the continuous parameters ${\mathbf{A}},{\mathbf{B}}$ into their discrete equivalents $\overline{{\mathbf{A}}},\overline{{\mathbf{B}}}$. This discretization is typically done through the Zero-Order Hold (ZOH) method given in Equation 2.
@@
While both S4 [17] and Mamba [16] utilize a similar discretization step as stated above in Equation 2, Mamba differentiates itself from S4 by conditioning the parameters ${\mathbf{\Delta}}\in\mathbb{R}^{B\times L\times D}$, ${\mathbf{B}}\in\mathbb{R}^{B\times L\times N}$ and ${\mathbf{C}}\in\mathbb{R}^{B\times L\times N}$, on the input $x\in\mathbb{R}^{B\times L\times D}$, through the S6 Selective Scan Mechanism, where $B$ is the batch size, $L$ is the sequence length, and $D$ is the feature dimension.
@@ 3.2 Overall Architecture
As shown in Figure 2 (a), our model uses a hierarchical architecture, similar to Swin Transformer [36], with four stages to efficiently process images at varying resolutions. Assuming an input image, $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, we first apply a Patch Embedding layer to divide the image into non-overlapping patches of size $4\times 4$ and embed each patch into a $C_{1}$-dimensional feature vector. The patch embedding layer is implemented using two $3\times 3$ convolutions with a stride of 2. This produces features maps of size $\frac{H}{4}\times\frac{W}{4}\times C_{1}$ at the first stage. These feature maps are passed to a stack of our Modulated Grouped Mamba blocks (as detailed in Section 3.3). In each subsequent stage, a down-sampling layer merges patches in a $2\times 2$ region, followed by another stack of our Modulated Grouped Mamba blocks. Hence, feature size at stages two, three and four are $\frac{H}{8}\times\frac{W}{8}\times C_{2}$, $\frac{H}{16}\times\frac{W}{16}\times C_{3}$, and $\frac{H}{32}\times\frac{W}{32}\times C_{4}$, respectively.
@@ 3.3 Modulated Group Mamba Layer
We present the overall operations of the proposed Modulated Group Mamba layer (Figure 2 (b)) for an input sequence $\mathbf{X}_{\textsf{in}}$, with dimensions $(B,H,W,C)$, where $B$ is the batch size, $C$ is the number of input channels and $H$/$W$ are the width and height of the feature map, in Equation 3.
@@
Here, $\mathbf{X}_{\textsf{GM}}$ is the output of Equation 6, $\mathbf{X}_{\textsf{CAM}}$ is the output of Equation 9, LN is the Layer Normalization [1] operation, FFN is the Feed-Forward Network as described by Equation 5, and $\mathbf{X}_{\textsf{out}}$ is the final output of the Modulated Group Mamba block. The individual operations, namely the GroupedMamba operator, the VSSS block used inside the GroupedMamba operator, and the CAM operator, are presented in Section 3.3.1, Section 3.3.2 and Section 3.3.3, respectively.
@@ 3.3.1 Visual Single Selective Scan (VSSS) Block
The VSSS block (Figure 2 (c)) is a token and channel mixer based on the Mamba operator. Mathematically, for an input token sequence $\mathbf{Z}_{\textsf{in}}$, the VSSS block performs the operations as described in Equation 4.
@@
Where $\mathbf{Z}_{\textsf{out}}$ is the output sequence, Mamba is the discretized version of the Mamba SSM operator as described in Equation 2.
@@ 3.3.2 Grouped Mamba Operator
Considering the motivation presented earlier in Section 3, we aim to design a variant of the Mamba [16] that is both computationally efficient and can effectively model the spatial dependencies of the input sequence. Given that Mamba is computationally inefficient on large number of channels $C$ in the input sequence, we propose a grouped variant of the operator, inspired by Grouped Convolutions. The Grouped Mamba operation is a variant of the VSSS block presented in Section 3.3.1, where the input channels are divided into groups, and the VSSS operator is applied separately to each group. Specifically, we divide the input channels into four groups, each of size $\frac{C}{4}$, and an independent VSSS block is applied to each group. To better model spatial dependencies in the input, each of the four groups scans in one of four directions across the token sequence: left-to-right, right-to-left, top-to-bottom, and bottom-to-top, as outlined in Figure 2 (d).
@@
Given the above definitions, the overall relation for the Grouped Mamba operator can be written as shown in Equation 6.
$\mathbf{X}_{\textsf{GM}}=\textsf{GroupedMamba}(\mathbf{X}_{\textsf{in}},\Theta)=\textsf{Concat}(&\textsf{VSSS}(\mathbf{X}_{\textsf{LR}},\Theta_{\textsf{LR}}),\\ &\textsf{VSSS}(\mathbf{X}_{\textsf{RL}},\Theta_{\textsf{RL}}),\\ &\textsf{VSSS}(\mathbf{X}_{\textsf{TB}},\Theta_{\textsf{TB}}),\\ &\textsf{VSSS}(\mathbf{X}_{\textsf{BT}},\Theta_{\textsf{BT}}))\end{aligned}$ (6)
@@ 3.3.3 Channel Affinity Modulation (CAM)
On its own, the Grouped Mamba operator may have a disadvantage in the form of limited information exchange across channels, given the fact that each operator in the group only operates over $\frac{C}{4}$ channels. To encourage the exchange of information across channels, we propose a Channel Affinity Modulation operator, which recalibrates channel-wise feature responses to enhance the representation power of the network. In this block, we first average pool the input to calculate the channel statistics as shown in Equation 7.
@@
where $\mathbf{X}_{\textsf{in}}$ is the input tensor, and AvgPool represents the global average pooling operation. Next comes the affinity calculation operation as shown in Equation 8.
@@
where $\delta$ and $\sigma$ represent non-linearity functions, and $W_{1}$ and $W_{2}$ are learnable weights. The role of $\sigma$ is to assign an importance weight to each channel to compute the affinity. The result of the affinity calculation is used to recalibrate the output of the Grouped Mamba operator, as shown in Equation 9.
@@
where $\mathbf{X}_{\textsf{CAM}}$ is the recalibrated output, $\mathbf{X}_{\textsf{GM}}$ is the concatenated output of the four VSSS groups from Equation 6, $\mathbf{X}_{\textsf{in}}$ is the input tensor, and $\textsf{Affinity}(\mathbf{X}_{\textsf{in}})$ are the channel-wise attention scores obtained from the channel affinity calculation operation in Equation 8.
@@ 3.4 Distilled Loss Function
As mentioned earlier in the motivation in Section 3, the Mamba training is unstable when scaled to large models [45]. To mitigate this issue, we propose to utilize a distillation objection alongside the standard cross-entropy objective. Knowledge distillation involves training a student model to learn from a teacher model’s behavior by minimizing a combination of the classification loss and distillation loss. The distillation loss is computed using the cross-entropy objective between the logits of the teacher and student models. Given the logits ($Z_{s}$) from the student model, logits ($Z_{t}$) from a teacher model (RegNetY-16G [47] in our case), the ground truth label $y$, and the hard decision of the teacher $y_{\mathrm{t}}=\mathrm{argmax}_{c}Z_{\mathrm{t}}(c)$, the joint loss function is defined as shown in Equation 10.
@@
where $\mathcal{L}_{\mathrm{CE}}$ is the cross-entropy objective and $\alpha$ is the weighting parameter. We experimentally show in Section 4 that training with this distillation objective stabilizes training, leading to consistent performance gains on larger model variants.

### v2

# GroupMamba: Efficient Group-Based Visual State Space Model

@@ Abstract
State-space models (SSMs) have recently shown promise in capturing long-range dependencies with subquadratic computational complexity, making them attractive for various applications. However, purely SSM-based models face critical challenges related to stability and achieving state-of-the-art performance in computer vision tasks. Our paper addresses the challenges of scaling SSM-based models for computer vision, particularly the instability and inefficiency of large model sizes. We introduce a parameter-efficient modulated group mamba layer that divides the input channels into four groups and applies our proposed SSM-based efficient Visual Single Selective Scanning (VSSS) block independently to each group, with each VSSS block scanning in one of the four spatial directions. The Modulated Group Mamba layer also wraps the four VSSS blocks into a channel modulation operator to improve cross-channel communication. Furthermore, we introduce a distillation-based training objective to stabilize the training of large models, leading to consistent performance gains. Our comprehensive experiments demonstrate the merits of the proposed contributions, leading to superior performance over existing methods for image classification on ImageNet-1K, object detection, instance segmentation on MS-COCO, and semantic segmentation on ADE20K. Our tiny variant with 23M parameters achieves state-of-the-art performance with a classification top-1 accuracy of 83.3% on ImageNet-1K, while being 26% efficient in terms of parameters, compared to the best existing Mamba design of same model size. Code and models are available at: https://github.com/Amshaker/GroupMamba
@@ 1 Introduction
Figure 1: Comparison in terms of Parameters vs. Top-1 Accuracy on ImageNet-1k [9]. Our GroupMamba-B achieves superior top-1 classification accuracy while reducing parameters by 36% compared to VMamba [35].
@@
However, Mamba, specifically the S6 algorithm, is known to be unstable for e.g., image classification, especially when scaled to large sizes [46]. Additionally, the Mamba model variant used in image classification, generally called the VSS (Visual State Space) block, can be more efficient in terms of parameters and compute requirements based on the number of channels. The VSS block includes extensive input and output projections along with depth-wise convolutions, whose parameters and compute complexities are directly proportional to the number of channels in the input. To address this issue, we propose a hierarchical-based Modulated Group Mamba layer that mitigates the aforementioned issues in a computation and parameter-efficient manner. The main contributions of our paper are:
@@
    To address the instability issue in the SSM-based architecture, we introduce a distillation-based training objective designed to stabilize models with a large number of parameters, leading to better performance and a smooth loss convergence trend.
@@
    We build a series of parameter-efficient generic classification models called “GroupMamba”, based on the proposed Modulated Group Mamba layer. Our tiny variant achieves 83.3% top-1 accuracy on ImageNet-1k [9] with $23M$ parameters and $4.5G$ FLOPs. Additionally, our base variant achieves top-1 accuracy of 84.5% with $57M$ parameters and $14G$ FLOPs, outperforming all recent SSM methods (see Fig. 1).
@@ 3 Method
  • Lack of Stability for Larger Models: We observe from [46] that Mamba [16] based image classification models with an MLP channel mixer are unstable when scaled to a large number of parameters. This instability can be seen in SiMBA-L (MLP) [46], which leads to sub-optimal classification results of $49\%$ accuracy. We mitigate this issue by introducing a Modulated Group Mamba design alongside a distillation objective (as presented in Sec. 3.4) that stabilizes the Mamba SSM training without modifying the channel mixer.
  • Efficient Improved Interaction: Given the computational impact of Mamba-based design on the number of channels, the proposed Modulated Group Mamba layer is computationally inexpensive and parameter efficient than the default Mamba and able to model both local and global information from the input tokens through multi-direction scanning. An additional Channel Affinity Modulation operator is proposed in this work to compensate for the limited channel interaction due to the grouped operation and enhance their interactions.
@@ 3.1 Preliminaries
State-Space Models: State-space models (SSMs) like S4 [17] and Mamba [16] are structured sequence architectures inspired by a combination of recurrent neural networks (RNNs) and convolutional neural networks (CNNs), with linear or near-linear scaling in sequence length. Derived from continuous systems, SSMs define and 1D function-to-function map for an input $x(t)\in\mathbb{R}^{L}\rightarrow y(t)\in\mathbb{R}^{L}$ via a hidden state $h(t)\in\mathbb{R}^{N}$. More formally, SSMs are described by the continuous time Ordinary Differential Equation (ODE) in Eq. 1.
@@
where $h(t)$ is the current hidden state, $h^{\prime}(t)$ is the updated hidden state, $x(t)$ is the current input, $y(t)$ is the output, ${\mathbf{A}}\in\mathbb{R}^{N\times N}$ is SSM’s evolution matrix, and ${\mathbf{B}}\in\mathbb{R}^{1\times N},{\mathbf{C}}\in\mathbb{R}^{1\times N}$ are the input and output projection matrices, respectively.
Discrete State-Space Models: To allow these models to be used in sequence modeling tasks in deep learning, they need to be discretized, converting the SSM from a continuous time function-to-function map into a discrete-time sequence-to-sequence map. S4 [17] and Mamba [16] are among the discrete adaptations of the continuous system, incorporating a timescale parameter ${\mathbf{\Delta}}$ to convert the continuous parameters ${\mathbf{A}},{\mathbf{B}}$ into their discrete equivalents $\overline{{\mathbf{A}}},\overline{{\mathbf{B}}}$. This discretization is typically done through the Zero-Order Hold (ZOH) method given in Eq. 2.
@@
While both S4 [17] and Mamba [16] utilize a similar discretization step as stated above in Eq. 2, Mamba differentiates itself from S4 by conditioning the parameters ${\mathbf{\Delta}}\in\mathbb{R}^{B\times L\times D}$, ${\mathbf{B}}\in\mathbb{R}^{B\times L\times N}$ and ${\mathbf{C}}\in\mathbb{R}^{B\times L\times N}$, on the input $x\in\mathbb{R}^{B\times L\times D}$, through the S6 Selective Scan Mechanism, where $B$ is the batch size, $L$ is the sequence length, and $D$ is the feature dimension.
@@ 3.2 Overall Architecture
As shown in Fig. 2 (a), our model uses a hierarchical architecture, similar to Swin-Transformer [36], with four stages to efficiently process images at varying resolutions. Assuming an input image, $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, we first apply a Patch Embedding layer to divide the image into non-overlapping patches of size $4\times 4$ and embed each patch into a $C_{1}$-dimensional feature vector. The patch embedding layer is implemented using two $3\times 3$ convolutions with a stride of 2. This produces features maps of size $\frac{H}{4}\times\frac{W}{4}\times C_{1}$ at the first stage. These feature maps are passed through $N_{1}$ blocks of our Modulated Grouped Mamba (as detailed in Sec. 3.3). In each subsequent stage, a down-sampling layer merges patches in a $2\times 2$ region, followed by another $N$ blocks of our Modulated Grouped Mamba layer. Hence, feature size at stages two, three and four are $\frac{H}{8}\times\frac{W}{8}\times C_{2}$, $\frac{H}{16}\times\frac{W}{16}\times C_{3}$, and $\frac{H}{32}\times\frac{W}{32}\times C_{4}$, respectively.
@@ 3.3 Modulated Group Mamba Layer
We present the overall operations of the proposed Modulated Group Mamba layer (Fig. 2 (b)) for an input sequence $\mathbf{X}_{\textsf{in}}$, with dimensions $(B,H,W,C)$, where $B$ is the batch size, $C$ is the number of input channels and $H$/$W$ are the width and height of the feature map, in Eq. 3.
@@
Here, $\mathbf{X}_{\textsf{GM}}$ is the output of Eq. 6, $\mathbf{X}_{\textsf{CAM}}$ is the output of Eq. 9, LN is the Layer Normalization [1] operation, FFN is the Feed-Forward Network as described by Eq. 5, and $\mathbf{X}_{\textsf{out}}$ is the final output of the Modulated Group Mamba block. The individual operations, namely the GroupedMamba operator, the VSSS block used inside the GroupedMamba operator, and the CAM operator, are presented in Sec. 3.3.1, Sec. 3.3.2 and Sec. 3.3.3, respectively.
@@ 3.3.1 Visual Single Selective Scan (VSSS) Block
The VSSS block (Fig. 2 (c)) is a token and channel mixer based on the Mamba operator, comprising of a Mamba block followed by a Feed-Forward Network, each with a LayerNorm before it. Mathematically, for an input token sequence $\mathbf{Z}_{\textsf{in}}$, the VSSS block performs the operations as described in Eq. 4.
@@
Where $\mathbf{Z}_{\textsf{out}}$ is the output sequence, Mamba is the discretized Mamba SSM operator as described in Eq. 2.
@@ 3.3.2 Grouped Mamba Operator
Considering the motivation presented earlier in Sec. 3, we aim to design a variant of the Mamba [16] that is both computationally efficient and can effectively model the spatial dependencies of the input sequence. Given that Mamba is computationally inefficient on large number of channels $C$ in the input sequence, we propose a grouped variant of the operator, inspired by Grouped Convolutions. The Grouped Mamba operation is a variant of the VSSS block presented in Sec. 3.3.1, where the input channels are divided into groups, and the VSSS operator is applied separately to each group. Specifically, we divide the input channels into four groups, each of size $\frac{C}{4}$, and an independent VSSS block is applied to each group. Hence, the proposed grouped mamba operator enhances the model efficiency by splitting channels into smaller groups. To better model spatial dependencies in the input, each of the four groups scans in one of four directions across the input: left-to-right, right-to-left, bottom-to-top, and top-to-bottom as outlined in Fig. 2 (d).
@@
Given the above definitions, the overall relation for the Grouped Mamba operator can be written as shown in Eq. 6.
$\mathbf{X}_{\textsf{GM}}=$ $(\displaystyle\textsf{GroupedMamba}(\mathbf{X}_{\textsf{in}},\Theta)=\textsf{Concat}\big{(}$6)
$\textsf{VSSS}(\mathbf{X}_{\textsf{LR}},\Theta_{\textsf{LR}}),\,\textsf{VSSS}(\mathbf{X}_{\textsf{RL}},\Theta_{\textsf{RL}}),$
$\textsf{VSSS}(\mathbf{X}_{\textsf{TB}},\Theta_{\textsf{TB}}),\,\textsf{VSSS}(\mathbf{X}_{\textsf{BT}},\Theta_{\textsf{BT}})\big{)}$
@@ 3.3.3 Channel Affinity Modulation (CAM)
On its own, the Grouped Mamba operator may have a disadvantage in the form of limited information exchange across channels, given the fact that each operator in the group only operates over $\frac{C}{4}$ channels. To encourage the exchange of information across channels, we propose a Channel Affinity Modulation operator, which recalibrates channel-wise feature responses to enhance the representation power of the network. In this block, we first average pool the input to calculate the channel statistics as shown in Eq. 7.
@@
where $\mathbf{X}_{\textsf{in}}$ is the input tensor, and AvgPool represents the global average pooling operation. Next comes the affinity calculation operation as shown in Eq. 8.
@@
where $\delta$ and $\sigma$ represent non-linearity functions, and $W_{1}$ and $W_{2}$ are learnable weights. The role of $\sigma$ is to assign an importance weight to each channel to compute the affinity. The result of the affinity calculation is used to recalibrate the output of the Grouped Mamba operator, as shown in Eq. 9.
@@
where $\mathbf{X}_{\textsf{CAM}}$ is the recalibrated output, $\mathbf{X}_{\textsf{GM}}$ is the concatenated output of the four VSSS groups from Eq. 6, $\mathbf{X}_{\textsf{in}}$ is the input tensor, and $\textsf{Affinity}(\mathbf{X}_{\textsf{in}})$ are the channel-wise attention scores obtained from the channel affinity calculation operation in Eq. 8.
While the average pooling and affinity procedure employed by the CAM module resembles the Squeeze-and-Excitation (SE) block [26], it introduces a distinct mechanism tailored explicitly for cross-channel attention within multi-group transformations. Specifically, CAM allows inter-group information exchange to overcome the inherent limitations of the ”Grouped Mamba Operator,” which inherently restricts interactions within individual groups. In contrast, SE blocks typically focus on recalibrating a single feature group and have not yet been investigated within the context of Mamba-based architectures.
@@ 3.4 Distilled Loss Function
As mentioned earlier in the motivation in Sec. 3, the Mamba training is unstable when scaled to large models [46]. To mitigate this issue, we propose to utilize a distillation objective alongside the standard cross-entropy objective. Knowledge distillation involves training a student model to learn from a teacher model’s behavior by minimizing a combination of the classification loss and distillation loss. The distillation loss is computed using the cross-entropy objective between the logits of the teacher and student models. Given the logits ($Z_{s}$) from the student model, logits ($Z_{t}$) from a teacher model (RegNetY-16G [48] in our case), the ground truth label $y$, and the hard decision of the teacher $y_{\mathrm{t}}=\mathrm{argmax}_{c}Z_{\mathrm{t}}(c)$, the joint loss function is defined as shown in Eq. 10.
@@
where $\mathcal{L}_{\mathrm{CE}}$ is the cross-entropy objective and $\alpha$ is the weighting parameter. We demonstrate in the supplementary material that incorporating the distilled loss enhances training stability, resulting in consistent performance improvements for larger model variants.